OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet? A. 4x-3 B. 2x^3+6x^2 C. 2x^2-6x D.2x^3-6x^2

OpenStudy (mathstudent55):

box is x feet wide W = x twice as long as it is wide If the width is x, then L = ?

OpenStudy (aric200):

The width of the box is x. The length of the box is = 2 * x. Height of the box = x - 3. Volume of box = length * width * height = 2x * x * (x-3) = 2x^2 * (x-3) Which= 2x^3 - 6x^2

OpenStudy (aric200):

Where are you confused?

OpenStudy (anonymous):

I understand now. I was just misreading. Thank you!